area of square field is 6400,how much long wire is required for boundry?
Answer by mananth(16946)

Area = 6400 sq.units
Area of square = L^2
L^2=6400
l= sqrt6400
L= 80 units
..
perimeter of square = 4*L
=4*80= 320 units